# Evaluation of the Pronunciation of Tajweed Rules Based on DNN as a Step Towards Interactive Recitation Learning

### v1

@@ Abstract
Proper recitation of the Quran, adhering to the rules of Tajweed, is crucial for preventing mistakes during recitation and requires significant effort to master. Traditional methods of teaching these rules are limited by the availability of qualified instructors and time constraints. Automatic evaluation of recitation can address these challenges by providing prompt feedback and supporting independent practice. This study focuses on developing a deep learning model to classify three Tajweed rules—separate stretching (Al Mad), tight noon (Ghunnah), and hide (Ikhfaa)—using the publicly available QDAT dataset, which contains over 1,500 audio recordings. The input data consisted of audio recordings from this dataset, transformed into normalized mel-spectrograms. For classification, the EfficientNet-B0 architecture was used, enhanced with a Squeeze-and-Excitation attention mechanism. The developed model achieved accuracy rates of 95.35%, 99.34%, and 97.01% for the respective rules. An analysis of the learning curves confirmed the model’s robustness and absence of overfitting. The proposed approach demonstrates high efficiency and paves the way for developing interactive educational systems for Tajweed study.
Keywords: Pronunciation evaluation, Tajweed rules, Mel-spectrogram, Transfer learning, DNN, CNN, Attention mechanism
@@ 3.3 Initial Approaches
3.4 Attention Mechanisms
To enhance the effectiveness of Tajweed rules classification, attention mechanisms were tested in addition to standard CNNs. These techniques help the model focus on the most significant features of the input data. In our case, attention mechanisms were applied to mel-spectrograms to recognize specific characteristics of Quranic recitation, such as vowel prolongation or nasal tones, which are difficult to capture with a standard CNN [13].
First, the Squeeze-and-Excitation (SE) mechanism was applied to the EfficientNet-B0 architecture. Two configurations were tested: 1) inserting the SE block before Global Average Pooling, and 2) placing it after pooling. In both cases, the attention mechanism enhanced the significant features of the data before passing them to the classifier. This led to an improvement in accuracy compared to the baseline model without attention. It is important to note that, although the SE mechanism is also used to reweight feature importance, it is not a classical attention mechanism. The SE block focuses on reweighting channels in convolutional layers rather than establishing dependencies between spatial positions.
@@ 3.5 Model Training
The model training was conducted to classify three Tajweed rules: separate stretching, tight noon, and hide. After testing various models and attention mechanisms, the approach combining EfficientNet-B0 with the Squeeze-and-Excitation (SE) mechanism after Global Average Pooling was selected. This choice allowed to achieve the best results on the QDAT dataset, which consists of 1505 audio recordings.
@@ 4 Result and Discussion
We trained the model using audio recordings from the QDAT dataset, converted into normalized mel-spectrograms. The classification was performed using the EfficientNet-B0 architecture, enhanced with the Squeeze-and-Excitation attention mechanism after Global Average Pooling.

### v3

@@ Abstract
Proper recitation of the Quran, adhering to the rules of Tajweed, is crucial for preventing mistakes during recitation and requires significant effort to master. Traditional methods of teaching these rules are limited by the availability of qualified instructors and time constraints. Automatic evaluation of recitation can address these challenges by providing prompt feedback and supporting independent practice. This study focuses on developing a deep learning model to classify three Tajweed rules—separate stretching (Al Mad), tight noon (Ghunnah), and hide (Ikhfaa)—using the publicly available QDAT dataset, which contains over 1,500 audio recordings. The input data consisted of audio recordings from this dataset, transformed into normalized mel-spectrograms. For classification, the EfficientNet-B0 architecture was used, enhanced with a Squeeze-and-Excitation block. The developed model achieved accuracy rates of 95.35%, 99.34%, and 97.01% for the respective rules. An analysis of the learning curves confirmed the model’s robustness and absence of overfitting. The proposed approach demonstrates high efficiency and paves the way for developing interactive educational systems for Tajweed study.
Keywords: Pronunciation evaluation, Tajweed rules, Mel-spectrogram, Transfer learning, DNN, CNN, Squeeze-and-Excitation block, Attention mechanism
@@ 3.3 Initial Approaches
3.4 Comparison of CNN Enhancement Methods
To enhance the effectiveness of Tajweed rules classification, attention mechanism and Squeeze-and-Excitation (SE) blocks were tested in addition to standard CNNs. These techniques help the model focus on the most significant features of the input data. While attention mechanisms dynamically identify key time-frequency regions in the mel-spectrogram, SE blocks adaptively rebalance channel strengths. In our case, these approaches were applied to mel-spectrograms to recognize specific characteristics of Quranic recitation, such as vowel prolongation or nasal tones, which are difficult to capture with a standard CNN [13].
First, the Squeeze-and-Excitation (SE) block was applied to the EfficientNet-B0 architecture. Two configurations were tested: 1) inserting the SE block before Global Average Pooling, and 2) placing it after pooling. In both cases, the SE block enhanced the significant features of the data before passing them to the classifier. This led to an improvement in accuracy compared to the baseline model without SE block.
@@ 3.5 Model Training
The model training was conducted to classify three Tajweed rules: separate stretching, tight noon, and hide. After testing various models and attention mechanisms, the approach combining EfficientNet-B0 with the SE block after Global Average Pooling was selected. This choice allowed to achieve the best results on the QDAT dataset, which consists of 1505 audio recordings.
@@ 4 Result and Discussion
We trained the model using audio recordings from the QDAT dataset, converted into normalized mel-spectrograms. The classification was performed using the EfficientNet-B0 architecture, enhanced with the Squeeze-and-Excitation block after Global Average Pooling.